To derive a relation between K_p and K_c for the reaction :

H₂(g) + CO₂(g) H₂O(g) + CO(g)


1
Expert's answer
2021-01-19T04:38:59-0500

H2(g)+C02(g)"\to" H20(g)+C0(g)

Kp=kc(RT)Δn

Kp= (PH2O) (PC0)/(PH2)(PCO2)

Kc=[H20][CO]/[H2][CO2]

PV=nRT

P=(n/V)RT

n/V=Concentration,[].

"\therefore" P=[]RT

Kp=[H20]RT[CO]RT/[H2]RT[CO2]RT

Kp=[H20][CO]/[H2][CO2]

"\therefore" Kp=Kc